OpenStudy (anonymous):

How do you solve an absolute value problem like: |-28.87|+2|19.3| ?

OpenStudy (anonymous):

you use the positive value

OpenStudy (anonymous):

28.87+2(19.3)= 28.87 + 38.6= 67.47

OpenStudy (anonymous):

generally, you take once the positive value, once the negative value, and according to the solutions you take intervals